Pearson Motors has a target capital structure of 45% debt and 55% common equity, with no preferred stock. The yield to maturity on the company's outstanding bonds is 11%, and its tax rate is 25%. Pearson's CFO estimates that the company's WACC is 10.60%. What is Pearson's cost of common equity? Do not round intermediate calculations. Round your answer to two decimal places.

  % ?
Answer #1

WACC is weithed Avg cost of Sources in capital Structure.

After Tax cost of Debt = YTM (1- Tax Rate )

= 11% * ( 1 - 0.25)

= 11% * 0.75

= 8.25%

WACC:

Source Weight Cost Wtd Cost
Equity 55% X 0.55X
Debt 45% 8.25% 0.0371
WACC 0.55X + 0.0371

Given WACC is 0.1060

Thus 055X + 0.0371 = 0.1060

0.55X = 0.1060 - 0.0371

= 0.0689

X = 0.0689 / 0.55

= 0.1253 i.e 12.53%

Cost of equity is 12.53%